OpenStudy (anonymous):

Determine whether the graphs of the pair of lines are parallel. x+12=y y-x=-8 The slope of the line x+12=y is?

OpenStudy (anonymous):

yes. both have the same slope

OpenStudy (anonymous):

soo what would the slope be? Im confused

OpenStudy (anonymous):

yeap both have same slope, so parallel. The slope of the given line is 1

OpenStudy (anonymous):

I can explain it to ya.

OpenStudy (anonymous):

just compare the given linear equation with y=mx+c where m is the slope and c is the y-intercept

OpenStudy (anonymous):

Please do

OpenStudy (anonymous):

x+12 =y that means y = x+12 the slope intercept form of a line is y = mx+b, where m is the slope y = (1)x+12 so slope is 1

OpenStudy (anonymous):

Ooooh , okay I got it now . Thanks

OpenStudy (anonymous):

The trick is to make both equations say y= blah blah If the number in front of the x are the same, then they're parallel. You know how to set them to where it says y=??????

OpenStudy (anonymous):

yes